OpenStudy (anonymous):

there are two small identical charges spheres each having charge Qand are seperated by a large distance. force on each sphere is F. now a third charge is brought in contact with any one sphere and kept at midpoint. calculate charge on this sphere in term of F..

OpenStudy (anonymous):

calculate charge on this middle sphere in term of F..

OpenStudy (anonymous):

if the 3rd particle is initially not charged and finally it is kept at mid point of distance of 2 particles,the charge on it is (R/2)*sqrt(4*pi*epsilon0*F).where R is the distance between two particles.

OpenStudy (anonymous):

no an is F ...

OpenStudy (anonymous):

how madam.......

OpenStudy (anonymous):

madam lol :P anyways ans is F in my notebook ...

OpenStudy (anonymous):

wanna see solution actaully i dont know how to do these kind qn thats why i posted it...

OpenStudy (experimentx):

is third sphere identical to two sphere??

OpenStudy (anonymous):

i guess we can assume as they didnt mention..

OpenStudy (experimentx):

My guess is the third charge shares equal amount of charge from contact ... i.e. if Q is the initial charge in one of spheres, then after contact Q/2 charge is remained while Q/2 is transferred to third charge ... and I think we have to calculate force on this sphere due to two other initial spheres.

OpenStudy (anonymous):

ok thnx and one more plz hold on posting it..

OpenStudy (experimentx):

\( Force = \frac{kQ(Q/2)}{(R/2)^2} + \frac{kQ(Q/2)}{(R/2)^2} = 4F ?? \),

OpenStudy (anonymous):

noo ans is F wait posting the solution now

OpenStudy (experimentx):

Oo ... is charge between two sphere specified?? I realized my mistake

OpenStudy (anonymous):

as i said nthng is given o i asume it +ve charge on both side

OpenStudy (experimentx):

 \[ Force = \frac{k (Q/2)^2}{(r/2)^2} - \frac{k (Q)(Q/2)}{(r/2)^2 } = -\frac{k (Q)^2}{(r)^2} = F \]